### v3

# QN-Mixer: A Quasi-Newton MLP-Mixer Model for Sparse-View CT Reconstruction

Abstract
Inverse problems span across diverse fields. In medical contexts, computed tomography (CT) plays a crucial role in reconstructing a patient’s internal structure, presenting challenges due to artifacts caused by inherently ill-posed inverse problems. Previous research advanced image quality via post-processing and deep unrolling algorithms but faces challenges, such as extended convergence times with ultra-sparse data. Despite enhancements, resulting images often show significant artifacts, limiting their effectiveness for real-world diagnostic applications. We aim to explore deep second-order unrolling algorithms for solving imaging inverse problems, emphasizing their faster convergence and lower time complexity compared to common first-order methods like gradient descent. In this paper, we introduce QN-Mixer, an algorithm based on the quasi-Newton approach. We use learned parameters through the BFGS algorithm and introduce Incept-Mixer, an efficient neural architecture that serves as a non-local regularization term, capturing long-range dependencies within images. To address the computational demands typically associated with quasi-Newton algorithms that require full Hessian matrix computations, we present a memory-efficient alternative. Our approach intelligently downsamples gradient information, significantly reducing computational requirements while maintaining performance. The approach is validated through experiments on the sparse-view CT problem, involving various datasets and scanning protocols, and is compared with post-processing and deep unrolling state-of-the-art approaches. Our method outperforms existing approaches and achieves state-of-the-art performance in terms of SSIM and PSNR, all while reducing the number of unrolling iterations required.
Figure 1: CT Reconstruction with 32 views of State-of-the-Art Methods. Comparative analysis with post-processing and first-order unrolling networks highlights QN-Mixer’s superiority in artifact removal, training time, and data efficiency.
[…]
1 Introduction
Computed tomography (CT) is a widely used imaging modality in medical diagnosis and treatment planning, delivering intricate anatomical details of the human body with precision. Despite its success, CT is associated with high radiation doses, which can increase the risk of cancer induction [50]. Adhering to the ALARA principle (As Low As Reasonably Achievable) [37], the medical community emphasizes minimizing radiation exposure to the lowest level necessary for accurate diagnosis. Numerous approaches have been proposed to reduce radiation doses while maintaining image quality. Among these, sparse-view CT emerges as a promising solution, effectively lowering radiation doses by subsampling the projection data, often referred to as the sinogram. Nonetheless, reconstructed images using the well-known Filtered Back Projection (FBP) algorithm [34], suffer from pronounced streaking artifacts (see Fig. 1), which can lead to misdiagnosis. The challenge of effectively reconstructing high-quality CT images from sparse-view data is gaining increasing attention in both the computer vision and medical imaging communities.
With the success of deep learning spanning diverse domains, initial image-domain techniques [6, 19, 59, 25, 28] have been introduced as post-processing tasks on the FBP reconstructed images, exhibiting notable accomplishments in artifact removal and structure preservation. However, the inherent limitations of these methods arise from their constrained receptive fields, leading to challenges in effectively capturing global information and, consequently, sub-optimal results.
To address this limitation, recent advances have seen a shift toward a dual-domain approach [29, 18, 49, 27], where post-processing methods turn to the sinogram domain. In this dual-domain paradigm, deep neural networks are employed to perform interpolation tasks on the sinogram data [24, 15], facilitating more accurate image reconstruction. Despite the significant achievements of post-processing and dual-domain methods, they confront issues of interpretability and performance limitations, especially when working with small datasets and ultra-sparse-view data, as shown in Fig. 1. To tackle these challenges, deep unrolling networks have been introduced [1, 16, 7, 51, 8, 11, 20, 54]. Unrolling networks treat the sparse-view CT reconstruction problem as an optimization task, resulting in a first-order iterative algorithm like gradient descent, which is subsequently unrolled into a deep recurrent neural network in order to learn the optimization parameters and the regularization term. Like post-processing techniques, unrolling networks have been extended to the sinogram domain [52, 56] to perform interpolation task.
Unrolling networks, as referenced in [44, 12, 36], exhibit remarkable performance across diverse domains. However, they suffer from slow convergence and high computational costs, as illustrated in Fig. 1, necessitating the development of more efficient alternatives [14]. More specifically, they confront two main issues: Firstly, they frequently grapple with capturing long-range dependencies due to their dependence on locally-focused regularization terms using CNNs. This limitation results in suboptimal outcomes, particularly evident in tasks such as image reconstruction. Secondly, the escalating computational costs of unrolling methods align with the general trend of increased complexity in modern neural networks. This escalation not only amplifies the required number of iterations due to the algorithm’s iterative nature but also contributes to their high computational demand.
To tackle the aforementioned issues, we introduce a novel second-order unrolling network for sparse-view CT reconstruction. In particular, to enable the learnable regularization term to apprehend long-range interactions within the image, we propose a non-local regularization block termed Incept-Mixer. Drawing inspiration from the multi-layer perceptron mixer [46] and the inception architecture [45], it is created to combine the best features from both sides: capturing long-range interactions from the attention-like mechanism of MLP-Mixer and extracting local invariant features from the inception block. This block facilitates a more precise image reconstruction. Second, to cut down on the computational costs associated with unrolling networks, we propose to decrease the required iterations for convergence by employing second-order optimization methods such as [21, 30]. We introduce a novel unrolling framework named QN-Mixer. Our approach is based on the quasi-Newton method that approximate the Hessian matrix using the Broyden-Fletcher-Goldfarb-Shanno (BFGS) update [13, 10, 57]. Furthermore, we reduce memory usage by working on a projected gradient (latent gradient), preserving performance while reducing the computational cost tied to Hessian matrix approximation. This adaptation enables the construction of a deep unrolling network, showcasing superlinear convergence. Our contributions are summarized as follows:
  • We introduce a novel second-order unrolling network coined QN-Mixer where the Hessian matrix is approximated using a latent BFGS algorithm with a deep-net learned regularization term.
  • We propose Incept-Mixer, a neural architecture acting as a non-local regularization term. Incept-Mixer integrates deep features from inception blocks with MLP-Mixer, enhancing multi-scale information usage and capturing long-range dependencies.
  • We demonstrate the effectiveness of our proposed method when applied to the sparse-view CT reconstruction problem on an extensive set of experiments and datasets. We show that our method outperforms state-of-the-art methods in terms of quantitative metrics while requiring less iterations than first-order unrolling networks.
2 Related Works
In this section, we present prior work closely related to our paper. We begin by discussing the general framework for unrolling networks in Sec. 2.1, which is based on the gradient descent algorithm. Subsequently, in Sec. 2.2 and Sec. 2.3, we delve into state-of-the-art methods in post-processing and unrolling networks, respectively.
2.1 Background
Inverse Problem Formulation for CT.  Image reconstruction problem in CT can be mathematically formalized as the solution to a linear equation in the form of:
${\bm{y}}={\bm{A}}{\bm{x}},$ (1)
where ${\bm{x}}\in\mathbb{R}^{n}$ is the (unknown) object to reconstruct with $n=h\times w$, ${\bm{y}}\in\mathbb{R}^{m}$ is the data (i.e. sinogram), where $m=n_{v}\times n_{d}$, $n_{v}$ and $n_{d}$ denote the number of projection views and detectors, respectively. ${\bm{A}}\in\mathbb{R}^{n\times m}$ is the forward model (i.e. discrete Radon transform [40]). The goal of CT image reconstruction is to recover the (unknown) object, ${\bm{x}}$, from the observed data ${\bm{y}}$. As the problem is ill-posed due to the missing data, the linear system in Eq. 1 becomes underdetermined and may have infinite solutions. Hence, reconstructed images suffer from artifacts, blurring, and noise. To address this issue, iterative reconstruction algorithms are utilized to minimize a regularized objective function with a $L^{2}$ norm constraint:
$\hat{{\bm{x}}}=\underset{{\bm{x}}}{\operatorname{arg}\,\operatorname{min}}\;J({\bm{x}})=\frac{\lambda}{2}\left\lVert{\bm{A}}{\bm{x}}-{\bm{y}}\right\rVert^{2}_{2}+{\mathcal{R}}({\bm{x}}),$ (2)
where ${\mathcal{R}}({\bm{x}})$ is the regularization term, balanced with the weight $\lambda$. Those ill-posed problems were initially addressed using optimization techniques, such as the truncated singular value decomposition (SVD) algorithm [42], or iterative approaches like the algebraic reconstruction technique (ART) [4], simultaneous ART (SART) [2], conjugate gradient for least squares (CGLS) [22], and total generalized variation regularization (TGV) [43]. Additionally, techniques such as total variation [47] and Tikhonov regularization [9] can be employed to enhance reconstruction results.
Deep Unrolling Networks.  By assuming that the regularization term in Eq. 2 (i.e. ${\mathcal{R}}$) is differentiable and convex, a simple gradient descent scheme can be applied to solve the optimization problem:
$&{\bm{x}}_{t+1}={\bm{x}}_{t}-\alpha\nabla_{\bm{x}}J({\bm{x}}_{t}),\\ &\text{where}\quad\nabla_{\bm{x}}J({\bm{x}}_{t})=\lambda{\bm{A}}^{\dagger}\left({\bm{A}}{\bm{x}}_{t}-{\bm{y}}\right)+\nabla_{\bm{x}}{\mathcal{R}}({\bm{x}}_{t}).\end{aligned}$ (3)
Here, $\alpha$ represents the step size (i.e. search step), and ${\bm{A}}^{\dagger}$ is the pseudo-inverse of ${\bm{A}}$.
Previous research [53, 16] has emphasized the limitations of optimization algorithms, such as the manual selection of the regularization term and the optimization hyper-parameters, which can negatively impact their performance, limiting their clinical application. Recent advancements in deep learning techniques have enabled automated parameter selection directly from the data, as demonstrated in [33, 7, 56, 11, 23, 38]. By allowing the terms in Eq. 3 to be dependent on the iteration, the gradient descent iteration becomes:
${\bm{x}}_{t+1}={\bm{x}}_{t}-\lambda_{t}{\bm{A}}^{\dagger}\left({\bm{A}}{\bm{x}}_{t}-{\bm{y}}\right)+{\mathcal{G}}({\bm{x}}_{t}),$ (4)
where ${\mathcal{G}}$ is a learned mapping representing the gradient of the regularization term. It is worth noting that the step size $\alpha$ in Eq. 3 is omitted as it is redundant when considering the learned components of the regularization term. Finally, Eq. 4 is unrolled into a deep recurrent neural network in order to learn the optimization parameters.
2.2 Post-processing Methods
Figure 2: Overall structure of the proposed QN-Mixer for sparse-view CT reconstruction, unrolled from Algorithm 2. The method leverages the advantages of the quasi-Newton method for faster convergence while incorporating a latent BFGS update.
Recent advances in sparse-view CT reconstruction leverage two main categories of deep learning methods: post-processing and dual-domain approaches. Post-processing methods, including RedCNN [6], FBPConvNet [19], and DDNet [59], treat sparse-view reconstruction as a denoising step using FBP reconstructions as input. While effective in addressing artifacts and reducing noise, they often struggle with recovering global information from extremely sparse data. To overcome this limitation, dual-domain methods integrate sinograms into neural networks for an interpolation task, recovering missing data [24, 15]. Dual-domain methods, surpassing post-processing ones, combine information from both domains. DuDoNet [29], an initial dual-domain method, connects image and sinogram domains through a Radon inversion layer. Recent Transformer-based dual-domain methods, such as DuDoTrans [49] and DDPTransformer [27], aim to capture long-range dependencies in the sinogram domain, demonstrating superior performance to CNN-based methods.
Self-supervised learning.  SSL methods [17, 5, 58, 48, 26], have been applied for CT reconstruction. For instance, [5] proposed an equivariant imaging paradigm through a training strategy that enforces measurement consistency and equivariance conditions. To ensure equitable comparisons, we focus on supervised methods in this work.
2.3 Advancements in Deep Unrolling Networks
Unrolling networks constitute a line of work inspired by popular optimization algorithms used to solve Eq. 2. Leveraging the iterative nature of optimization algorithms, as presented in Eq. 4, unrolling networks aim to directly learn optimization parameters from data. These methods have found success in various inverse problems, including sparse-view CT [7, 20, 52, 56, 54], limited-angle CT [11, 51, 8], low-dose CT [1, 16], and compressed sensing MRI [44, 12].
First-order.  One pioneering unrolling network, Learned Primal-Dual reconstruction [1], replaces traditional proximal operators with CNNs. In contrast, LEARN [7] and LEARN++ [56] directly unroll the optimization algorithm from Eq. 4 into a deep recurrent neural network. More recently, Transformers [3, 31] have been introduced into unrolling networks, such as RegFormer [54] and HUMUS-Net [12]. While achieving commendable performance, these methods require more computational resources than traditional CNN-based unrolling networks and incur a significant memory footprint due to linear scaling with the number of unrolling iterations.
Second-order.  To address this, a new category of unrolling optimization methods has emerged [14], leveraging second-order techniques like the quasi-Newton method [21, 13, 10]. These methods converge faster, reducing computational demands, but struggle with increased memory usage due to Hessian matrix approximation and their application is limited to small-scale problems [30, 57]. In contrast our method propose a memory-efficient approach by operating within the latent space of gradient information (i.e. $\nabla_{\bm{x}}J({\bm{x}})$ in Eq. 3).
Data: ${\bm{y}}$ (sparse sinogram)
Manual choice of the regularization term ${\mathcal{R}}$;
${\bm{H}}_{0}\leftarrow{\bm{I}}^{n\times n}$;
${\bm{x}}_{0}\leftarrow{\bm{A}}^{\dagger}{\bm{y}}$;
for $t\in\{0,\dots,T-1\}$ do
       ${\bm{s}}_{t}\leftarrow-{\bm{H}}_{t}\nabla_{{\bm{x}}}J({\bm{x}}_{t})$
      ${\bm{x}}_{t+1}\leftarrow{\bm{x}}_{t}+{\bm{s}}_{t}$
      ${\bm{z}}_{t}\leftarrow\nabla_{{\bm{x}}}J({\bm{x}}_{t+1})-\nabla_{{\bm{x}}}J({\bm{x}}_{t})$
      $\rho_{t}\leftarrow 1/({\bm{z}}_{t}^{\mathsf{T}}{\bm{s}}_{t})$
      ${\bm{H}}_{t+1}\leftarrow({\bm{I}}-\rho_{t}{\bm{s}}_{t}{\bm{z}}_{t}^{\mathsf{T}}){\bm{H}}_{t}({\bm{I}}-\rho_{t}{\bm{z}}_{t}{\bm{s}}_{t}^{\mathsf{T}})+\rho_{t}{\bm{s}}_{t}{\bm{s}}_{t}^{\mathsf{T}}$
Algorithm 1 Quasi-Newton for sparse-view CT
3 Methodology
QN-Mixer is a novel second-order unrolling network inspired by the quasi-Newton (Sec. 3.1) method. It approximates the inverse Hessian matrix with a latent BFGS algorithm and includes a non-local regularization term, Incept-Mixer, designed to capture non-local relationships (Sec. 3.2). To cope with the significant computational burden associated with the full approximation of the inverse Hessian matrix, we use a latent BFGS algorithm (Sec. 3.3). An overview of the proposed method is depicted in Fig. 2, and the complete algorithm is presented in Sec. 3.4.
3.1 Quasi-Newton method
The quasi-Newton method can be applied to solve Eq. 2 and the iterative optimization solution is expressed as:
${\bm{x}}_{t+1}={\bm{x}}_{t}-\alpha_{t}{\bm{H}}_{t}\nabla_{\bm{x}}J({\bm{x}}_{t}),$ (5)
where ${\bm{H}}_{t}\in\mathbb{R}^{n\times n}$ represents the inverse Hessian matrix approximation at iteration $t$, and $\alpha_{t}$ is the step size. The BFGS method updates the Hessian matrix approximation in each iteration. This matrix is crucial for understanding the curvature of the objective function around the current point, guiding us to take more efficient steps and avoiding unnecessary zigzagging. In the classical BFGS approach, the line search adheres to Wolfe conditions [13, 10]. A step size of $\alpha_{t}=1$ is attempted first, ensuring eventual acceptance for superlinear convergence [21]. In our approach, we adopt a fixed step size of $\alpha_{t}=1$. The algorithm is illustrated in Algorithm 1.
3.2 Regularization term: Incept-Mixer
Figure 3: Architecture of our regularization block. It is referred to as “Incept-Mixer" and denoted as ${\mathcal{G}}$ in Eq. 4
Recent research on unrolling networks has often focused on selecting the representation of the regularization term gradient (i.e. ${\mathcal{G}}$ in Eq. 4), ranging from conv-nets [7, 56, 44] to more recent attention-based nets [54, 12]. In alignment with this trend, we introduce a non-local regularization block named Incept-Mixer and depicted in, Fig. 3. This block is crafted by drawing inspiration from both the multi-layer perceptron mixer [46] and the inception architecture [45], leveraging the strengths of each: capturing long-range interactions through the attention-like mechanism of MLP-Mixer and extracting local invariant features from the inception block. This design choice is evident in the ablation study (see Tab. 6) where Incept-Mixer outperforms both alternatives.
Starting from an image ${\bm{x}}_{t}\in\mathbb{R}^{h\times w\times c}$ at iteration $t$, we pass it through an Inception block to create a feature map ${\bm{f}}_{t}\in\mathbb{R}^{h\times w\times d}$, where $d$ is the depth of features. Subsequently, ${\bm{f}}_{t}$ undergoes patchification using a CNN with a kernel size and stride of $p$, representing the patch size. This process yields patch embeddings, ${\bm{e}}_{t}=\mathrm{patchify}({\bm{f}}_{t})\in\mathbb{R}^{\frac{h}{p}\times\frac{w}{p}\times d}$. These embeddings are then processed through a Mixer Layer with token and channel MLPs, layer normalization, and skip connections for inter-layer information flow, following [46]:
$\mathrm{Mixer}({\bm{e}}_{t})=\mathrm{Mix}(\mathrm{MLP}_{\texttt{c}},\mathrm{Mix}(\left[\mathrm{MLP}_{\texttt{h}},\mathrm{MLP}_{\texttt{w}}\right],{\bm{e}}_{t}),$ (6)
where $\mathrm{Mix}(\mathrm{Layer},{\bm{e}}_{t})=\mathrm{Layer}(\mathrm{LN}({\bm{e}}_{t}))+{\bm{e}}_{t}$, with $\mathrm{LN}$ as layer normalization. $\mathrm{MLP}_{\texttt{h}}$, $\mathrm{MLP}_{\texttt{w}}$ are applied to height and width features, respectively, and $\mathrm{MLP}_{\texttt{c}}$ to rows and shared. Finally, after $N$ such mixer layers, the regularized sample is transformed back to an image through a patch expansion step to obtain ${\mathcal{G}}({\bm{x}}_{t})$. Consequently, the iterative optimization solution is as follows:
$&{\bm{x}}_{t+1}={\bm{x}}_{t}-{\bm{H}}_{t}\nabla_{\bm{x}}J({\bm{x}}_{t}),\\ &\text{where}\quad\nabla_{\bm{x}}J({\bm{x}}_{t})=\lambda_{t}{\bm{A}}^{\dagger}\left({\bm{A}}{\bm{x}}_{t}-{\bm{y}}\right)+{\mathcal{G}}({\bm{x}}_{t}).\end{aligned}$ (7)
Here, ${\mathcal{G}}({\bm{x}}_{t})$ denotes the Incept-Mixer model, representing the learned gradient of the regularization term.
3.3 Latent BFGS update
We propose a memory-efficient latent BFGS update. Drawing inspiration from LDMs [41], at step $t$, given the gradient value $\nabla_{{\bm{x}}}J({\bm{x}}_{t})\in\mathbb{R}^{h\times w\times c}$, the encoder $\mathcal{E}$ encodes it into a latent representation ${\bm{r}}_{t}=\mathcal{E}(\nabla_{{\bm{x}}}J({\bm{x}}_{t}))\in\mathbb{R}^{l_{h}\cdot l_{w}}$. Importantly, the encoder downsamples the gradient by a factor ${\bm{f}}_{\mathcal{E}}=\frac{h}{h_{l}}=\frac{w}{w_{l}}$. Throughout the paper, we explore different downsampling factors (see Tab. 5) ${\bm{f}}_{\mathcal{E}}=2^{k}$, where $k\in{\mathbb{N}}$ is the number of downsampling stacks. Encoding the gradient reduces the optimization variable size of BFGS (i.e. ${\bm{H}}_{t}\in\mathbb{R}^{(l_{h}\cdot l_{w})\times(l_{h}\cdot l_{w})}$), thereby decreasing the computational cost associated with high memory demand. The direction is then computed in the latent space $s_{t}=-{\bm{H}}_{t}{\bm{r}}_{t}$, and finally, the decoder $\mathcal{D}$ reconstructs the update from the latent direction, giving $\mathcal{D}(s_{t})=\mathcal{D}(-{\bm{H}}_{t}\mathcal{E}(\nabla_{{\bm{x}}}J({\bm{x}}_{t})))\in\mathbb{R}^{h\times w\times c}$. It is noteworthy that $\mathcal{E}$ and $\mathcal{D}$ are shared across the algorithm iterations, as shown in Fig. 2.
3.4 Proposed algorithm of QN-Mixer
Data: ${\bm{y}}$ (sparse sinogram)
${\bm{H}}_{0}\leftarrow{\bm{I}}^{(l_{h}\cdot l_{w})\times(l_{h}\cdot l_{w})}$;
${\bm{x}}_{0}\leftarrow{\bm{A}}^{\dagger}{\bm{y}}$;
${\bm{r}}_{0}\leftarrow\mathcal{E}(\nabla_{{\bm{x}}}J({\bm{x}}_{0}))$;
for $t\in\{0,\dots,T-1\}$ do
       ${\bm{s}}_{t}\leftarrow-{\bm{H}}_{t}{\bm{r}}_{t}$
      ${\bm{x}}_{t+1}\leftarrow{\bm{x}}_{t}+\mathcal{D}({\bm{s}}_{t})$
      ${\bm{r}}_{t+1}\leftarrow\mathcal{E}(\nabla_{{\bm{x}}}J({\bm{x}}_{t+1}))$
      ${\bm{z}}_{t}\leftarrow{\bm{r}}_{t+1}-{\bm{r}}_{t}$
      $\rho_{t}\leftarrow 1/({\bm{z}}_{t}^{\mathsf{T}}{\bm{s}}_{t})$
      ${\bm{H}}_{t+1}\leftarrow({\bm{I}}-\rho_{t}{\bm{s}}_{t}{\bm{z}}_{t}^{\mathsf{T}}){\bm{H}}_{t}({\bm{I}}-\rho_{t}{\bm{z}}_{t}{\bm{s}}_{t}^{\mathsf{T}})+\rho_{t}{\bm{s}}_{t}{\bm{s}}_{t}^{\mathsf{T}}$
Algorithm 2 QN-Mixer (latent BFGS update)
Our method, builds on the BFGS update [13, 10] rank-one approximation for the inverse Hessian. This approximation serves as a preconditioning matrix, guiding the descent direction. In contrast to [14], which directly learns the inverse Hessian approximation from data, our approach incorporates the mathematical equations of the BFGS algorithm for more accurate approximations. The full QN-Mixer algorithm is illustrated in Algorithm 2.
Method No noise $(N_{0}=0)$ Low noise $(N_{1}=10^{6})$ High noise $(N_{2}=5\times 10^{5})$
$n_{v}=32$ $n_{v}=64$ $n_{v}=128$ $n_{v}=32$ $n_{v}=64$ $n_{v}=128$ $n_{v}=32$ $n_{v}=64$ $n_{v}=128$
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
FBP 22.65 40.49 27.29 57.94 33.04 79.50 22.09 32.73 26.51 49.56 31.69 71.09 19.05 15.56 22.71 25.74 26.52 40.87
FBPConvNet [19] 30.32 85.11 35.42 90.15 39.71 94.64 30.20 84.46 35.09 89.72 39.06 94.08 29.91 82.52 34.13 87.85 36.89 91.28
DuDoTrans [49] 30.48 84.70 35.37 91.87 40.62 96.41 30.34 83.72 35.36 91.42 39.75 95.49 30.09 81.83 34.09 88.67 37.08 93.44
Learned PD [1] 35.88 92.09 41.03 96.28 43.33 97.31 35.78 92.21 39.03 94.79 41.65 96.44 33.80 89.23 37.34 93.23 39.17 94.69
LEARN [7] 37.58 94.65 42.26 97.25 43.11 97.57 36.95 93.63 39.91 95.82 42.17 97.11 34.38 90.51 37.15 93.53 39.38 95.18
RegFormer [54] 38.71 95.42 43.56 97.76 47.95 98.98 37.21 94.73 41.65 96.92 44.38 98.02 35.93 92.78 38.53 94.84 40.52 96.19
QN-Mixer (ours) 39.51 96.11 45.57 98.48 50.09 99.32 37.50 94.92 42.46 97.70 44.27 98.11 35.91 92.49 38.73 94.92 40.51 96.27
Table 1: Quantitative evaluation on AAPM of state-of-the-art methods (PSNR in dB and SSIM in %). Bold: Best, under: second best.
Figure 4: Visual comparison on AAPM. From top to bottom: the results under the following conditions: first $(n_{v}{=}32,N_{1})$, second $(n_{v}{=}64,N_{1})$, third $(n_{v}{=}32,N_{0})$. The last row presents out-of-distribution (OOD) results with a randomly overlaid circle on a test image. The display window is set to $\left[-1000,800\right]$ HU.
4 Experiments
In this section, we initially present our experimental settings, followed by a comparison of our approach with other state-of-the-art CT reconstruction methods. Finally, we delve into the contribution analysis of each component in our model.
Figure 5: Visual comparison on DeepLesion of state-of-the-art methods. Rows display results under different conditions: $(n_{v}{=}64,N_{1})$ and $(n_{v}{=}128,N_{1})$. Display windows are set to $\left[-1000,800\right]$ HU for the first row and $\left[-200,300\right]$ HU for the second row.
4.1 Experimental Setup
Datasets.  We evaluate our method on two widely used datasets: the “2016 NIH-AAPM-Mayo Clinic Low-Dose CT Grand Challenge” dataset (AAPM) [35] and the DeepLesion dataset [55]. The AAPM dataset comprises 2378 full-dose CT images from 10 patients, while DeepLesion is the largest publicly accessible multi-lesion real-world CT dataset, including 4427 unique patients.
Implementation details.  For AAPM, we select $1920$ training images from 8 patients, $244$ validation images from $1$ patient, and $214$ testing images from the last patient. For DeepLesion, we select a subset of $2000$ training images and $300$ testing images randomly from the official splits. All images are resized to $256\times 256$ pixels. To simulate the forward and backprojection operators, we use the Operator Discretization Library (ODL) [39] with a 2D fan-beam geometry ($512$ detector pixels, source-to-axis distance of $600$ mm, axis-to-detector distance of $290$ mm). Sparse-view CT images are generated with $n_{v}\in\{32,64,128\}$ projection views, uniformly sampled from a full set of $512$ views covering $\left[0,2\pi\right]$. To mimic real-world CT images, we introduce mixed noise to the sinograms, combining 5% Gaussian noise and Poisson noise with an intensity of $1\times 10^{6}$.
[…]
Evaluation metrics.  Following established evaluation protocols [1, 49, 54], we employ the structural similarity index measure (SSIM) with parameters: level $5$, a Gaussian kernel of size $11$, and standard deviation $1.5$, as our primary performance metric. Furthermore, we supplement our assessment with the peak signal-to-noise ratio (PSNR).
State-of-the-art baselines.  We compare QN-Mixer to multiple state-of-the-art competitors: (1) post-processing based denoising methods, i.e., FBPConvNet [19], and DuDoTrans [49]; (2) first-order unrolling reconstruction networks, i.e., Learned Primal-Dual [1], LEARN [7], and RegFormer [54]. Note that we replace the pseudo-inverse operator used by LEARN with the FBP algorithm, as it has been demonstrated to be more effective according to [54]. To ensure a fair comparison, we utilize the code-base released by the authors when possible or meticulously implement the methods based on the details provided in their papers. All approaches undergo training and testing on the same datasets, as elaborated in implementation details.
4.2 Comparison with state-of-the-art methods
Method $n_{v}=32$ $n_{v}=64$ $n_{v}=128$
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
FBP 21.55 31.65 26.07 47.17 31.49 69.63
FBPConvNet [19] 30.74 80.41 34.64 87.36 38.69 92.94
DuDoTrans [49] 32.11 79.86 36.02 88.14 40.47 93.81
Learned PD [1] 34.02 88.44 37.56 92.46 40.79 95.32
LEARN [7] 35.76 92.12 39.83 95.66 41.34 96.21
RegFormer [54] 37.38 93.89 41.70 96.78 46.10 98.39
QN-Mixer (ours) 39.39 95.67 43.75 97.73 48.62 98.64
Table 2: Quantitative evaluation on DeepLesion for state-of-the-art methods (PSNR in dB and SSIM in %). With Poisson noise level of $N_{1}=10^{6}$. Bold: Best, under: second best.
Quantitative comparison.  We compared our model with state-of-the-art baselines on two public datasets. For AAPM, models were trained and tested across three projection views ($n_{v}\in\{32,64,128\}$) and three noise levels, namely no noise $N_{0}=0$, low noise $N_{1}=10^{6}$, and high noise $N_{2}=5\times 10^{5}$ (see Tab. 1). For DeepLesion, models were trained and tested on the same three projection views and a noise level of $N_{1}=10^{6}$ (see Tab. 2). Visual results are provided in Fig. 4 (AAPM) and Fig. 5 (DeepLesion). Impressively, our method achieves state-of-the-art results on DeepLesion across all projection views. It outperforms the second-best baseline, RegFormer, with an average improvements of $+2.23$ dB in PSNR and $+1.02$% in SSIM. On AAPM without noise, we achieve state-of-the-art results across all projection views and improve the second best by an average $+1.65$ dB and $+0.58\%$. In the presence of low noise, QN-Mixer achieves state-of-the-art results performance in all cases except $n_{v}{=}128$ with $-0.11$ dB and shows an average improvements of $+0.33$ dB and $+0.35$% over RegFormer. With high noise, our method performs nearly on par in $n_{v}{=}32$ ($-0.02$ dB and $-0.29$%), achieves state-of-the-art in $n_{v}{=}64$ ($+0.2$ dB and $+0.08$%), and competes closely in $n_{v}{=}128$ ($-0.01$ dB and $+0.08$%). As noise increases, we attribute the decline in improvement to the compressed gradient information in the latent BFGS, influenced by sinogram changes, and the utilization of the FBP algorithm instead of the pseudo-inverse.
Performance comparison on OOD textures.
Method $n_{v}=32$ $n_{v}=64$ $n_{v}=128$
PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
FBP 21.38 33.36 26.08 50.29 31.43 73.06
FBPConvNet [19] 28.05 75.96 32.50 82.90 35.45 88.14
DuDoTrans [49] 28.11 68.17 32.71 83.26 36.41 90.36
Learned PD [1] 31.96 87.10 36.40 92.57 37.63 93.17
LEARN [7] 34.48 90.15 36.89 91.85 38.32 94.67
RegFormer [54] 34.49 89.98 36.95 91.48 38.02 92.44
QN-Mixer (ours) 36.84 94.84 42.11 97.78 45.69 98.82
Table 3: Quantitative evaluation on out-of-distribution (OOD) AAPM test dataset of state-of-the-art methods (PSNR in dB and SSIM in %). Bold: Best, under: second best.
We evaluate frozen model performance on CT images featuring a randomly positioned white circle with noise-free sinograms, as illustrated in the third row of Fig. 4. The rationale and details are provided in the supplementary material. In Tab. 3, QN-Mixer attains state-of-the-art results across all $n_{v}$ views. First-order unrolling networks such as LEARN and RegFormer exhibit significant PSNR degradation of $-3.1$ dB and $-4.22$ dB, respectively, for $n_{v}{=}32$, while our method demonstrates a milder degradation of $-2.67$ dB.
Visual comparison.  As it can be seen on Fig. 4 and Fig. 5, FBPConvNet and DuDoTrans show significant blurring and pronounced artifacts when $n_{v}{=}32$. While Learned PD and LEARN show satisfactory performance, they struggle with intricate details, like in the liver and spine. In contrast, RegFormer produces high-quality images but faces challenges in generalizing to OOD data. QN-Mixer excels in producing high-quality images with fine details, even under challenging conditions such as $n_{v}{=}32$ views and OOD data.
Method #Iters Epoch time (s) Time (ms) #Params (M) Memory (GB)
Post-processing based denoising
FBPConvNet [19] - 68 12.4 31.1 1.30
DuDoTrans [49] - 92 60.1 15.0 1.38
First-order unrolling reconstruction networks
Learned PD [1] 10 82 47.2 0.25 0.81
LEARN [7] 30 780 679.8 4.50 1.85
RegFormer [54] 18 700 598.9 5.00 10.19
Second-order unrolling Quasi-Newton
QN-Mixer (ours) 14 594 610.2 8.50 7.83
Table 4: Comparison of computational efficiency. Training epoch time is reported in seconds, #Params in M and memory costs for state-of-the-art methods on AAPM with $n_{v}=32$ views.
Efficiency comparison.  The results in Tab. 4 show that QN-Mixer is more computationally efficient than RegFormer, with a 1.3×1.3$\times$1.3 × reduction in memory usage. Furthermore, our training time demonstrates a significant enhancement, realizing a speed improvement of $106$ seconds per epoch compared to first-order unrolling methods like LEARN and RegFormer. Additionally, our method requires only $14$ iterations, in contrast to the $30$ and $18$ iterations needed by LEARN and RegFormer, respectively.
Hessian size PSNR $\uparrow$ SSIM $\uparrow$
$8^{2}\times 8^{2}$ 35.69 93.71
$16^{2}\times 16^{2}$ 38.11 95.31
$32^{2}\times 32^{2}$ 39.37 96.01
[…]
Table 5: Ablation on the inverse Hessian approximation size.
Method PSNR $\uparrow$ SSIM $\uparrow$
QN with different learned regularization
Inception 31.65 85.28
U-Net 34.29 92.92
MLP-Mixer 36.89 93.87
Incept-Mixer 39.51 96.11
Pseudo-inverse ${\bm{A}}^{\dagger}$ vs Filtered Back Projection (FBP)
QN-Mixer+${\bm{A}}^{\dagger}$ 38.94 95.83
QN-Mixer+FBP 39.51 96.11
First vs second order Quasi-Newton (QN)
Incept-Mixer+first-order 37.45 94.25
Incept-Mixer+QN 39.51 96.11
Table 6: QN-Mixer ablation.
4.3 Ablation Study
In this section, we leverage the AAPM dataset with $n_{v}{=}32$ views by default, and no noise is introduced to the sinogram.
Inverse Hessian approximation size.  The results in Tab. 5 emphasize the significant impact of the inverse Hessian approximation size on our performance. When too small, a notable degradation is observed (e.g., $8^{2}\times 8^{2}$), while larger sizes result in performance improvements as the approximation approaches the full inverse Hessian. Exceeding $64^{2}\times 64^{2}$ was unfeasible in our experiments due to memory constraints.
Figure 6: Ablation on the number of unrolling iterations. Left: PSNR (dB); Right: SSIM (%)
Number of unrolling iterations.  In Fig. 6, we visually depict the influence of the number of unrolling iterations on the performance of QN-Mixer and RegFormer. Notably, the performance of both methods shows improvement with an increase in the number of iterations. When subjected to an equal number of iterations, our method consistently surpasses RegFormer in performance. Remarkably, we achieve comparable results to RegFormer even with only 10 iterations, demonstrating the efficiency of our approach.
Regularization term.  In Tab. 6, we evaluate the impact of the regularization term in our framework. Our Incept-Mixer is compared against various learned alternatives, including the Inception block [45] and MLP-Mixer block [46]. Additionally, employing the pseudo-inverse ${\bm{A}}^{\dagger}$ instead of the FBP results in a less pronounced degradation ($-0.57$ dB and $-0.28$%), enhancing the interpretability of QN-Mixer. Finally, we test our Incept-Mixer in the first-order framework, highlighting the significance of the second-order latent BFGS approximation with a significant improvement ($+2.06$ dB and $+1.86\%$).
5 Conclusion
In this paper, we investigate the application of deep second-order unrolling networks for tackling imaging inverse problems. To this end, we introduce QN-Mixer, a quasi-Newton inspired algorithm where a latent BFGS method approximates the inverse Hessian, and our Incept-Mixer serves as the non-local learnable regularization term. Extensive experiments confirm the successful sparse-view CT reconstruction by our model, showcasing superior performance with fewer iterations than state of-the-art methods. In summary, this research offers a fresh perspective that can be applied to any iterative reconstruction algorithm. A limitation of our work is the memory requirements associated with quasi-Newton algorithm. We introduced a memory efficient alternative by projecting the gradient to a lower dimension, successfully addressing the CT reconstruction problem. However, its applicability to other inverse problems may be limited. In future work, we aim to extend our approach to handle larger Hessian sizes, broadening its application to a range of problems.
6 Acknowledgments
This work was granted access to the HPC resources of IDRIS under the allocation 2021-[AD011012741] provided by GENCI and supported by DIM Math Innov funding.
Supplementary Material
[…]
1
Appendix A Inverse Hessian approximation
Figure 7: Visualization of the inverse Hessian approximation across iterations. Observe the subtle changes between each iteration, attributed to the influence of the objective function used to estimate $H_{t}$ (see Eq. 10).
Figure 8: Inverse Hessian approximation rows visualization. We present the first $40$ rows for the $9^{th}$ and $14^{th}$ inverse Hessian approximations on the first and second lines, respectively. Each row is of size $64^{2}$, reshaped into a $64\times 64$ image. The corresponding image reconstructions are shown on the left, along with PSNR (dB) and SSIM (%) values at the top.
A.1 Approximation via optimization
The fundamental idea is to iteratively build a recursive approximation by utilizing curvature information along the trajectory. It is crucial to emphasize that a quadratic approximation offers a direction that can be leveraged within the iterative update scheme. This direction is defined by the equation:
${\bm{x}}_{t+1}={\bm{x}}_{t}+\alpha_{t}{\bm{d}}_{t}.$ (8)
In order to determine the direction ${\bm{d}}_{t}$, we can employ a quadratic approximation of the objective function. This approximation can be expressed as:
$J({\bm{x}}_{t}+{\bm{d}})\approx m_{t}({\bm{d}})=J({\bm{x}}_{t})+\nabla J({\bm{x}}_{t})^{\mathsf{T}}{\bm{d}}+\frac{1}{2}{\bm{d}}^{\mathsf{T}}{\bm{B}}_{t}{\bm{d}},$ (9)
where $J({\bm{x}}_{t})$ represents the objective function evaluated at the current point ${\bm{x}}_{t}$, $\nabla J({\bm{x}}_{t})$ denotes the gradient of the objective function at ${\bm{x}}_{t}$, ${\bm{B}}_{t}\in\mathbb{R}^{n\times n}$ corresponds to the approximation of the Hessian matrix. By minimizing the right-hand side of the quadratic approximation in Eq. 9, we can determine the optimal direction ${\bm{d}}_{t}$. Taking the derivative of $m_{t}({\bm{d}})$ with respect to ${\bm{d}}$ and setting it to zero, we obtain:
$\frac{\nabla m_{t}({\bm{d}})}{\nabla{\bm{d}}}={\bm{d}}_{t}{\bm{B}}_{t}+\nabla J({\bm{x}}_{t})\xrightarrow{\nabla m_{t}({\bm{d}})=0}{\bm{d}}_{t}=-{\bm{B}}_{t}^{-1}\nabla J({\bm{x}}_{t}),$
by substituting this result in Eq. 8 we obtain:
${\bm{x}}_{t+1}={\bm{x}}_{t}-\alpha_{t}{\bm{B}}_{t}^{-1}\nabla J({\bm{x}}_{t}).$
The objective is to ensure that the curvature along the trajectory is consistent. In other words, at the last two iterations, $m_{t+1}$ should match the gradient $\nabla J({\bm{x}}_{t})$ in the following way:
$\nabla m_{t+1}\bigg{|}_{{\bm{d}}=0}=\nabla J({\bm{x}}_{t+1}),\ \nabla m_{t+1}\bigg{|}_{{\bm{d}}=-\alpha_{t}{\bm{d}}_{t}}=\nabla J({\bm{x}}_{t}).$
This condition ensures that the quadratic approximation captures the correct curvature information along the trajectory, allowing for accurate optimization and convergence of the algorithm. By evaluating $\nabla m_{t+1}(\cdot)$ at the point $-\alpha_{t}{\bm{d}}_{t}$ we obtain:
$\alpha_{t}{\bm{B}}_{t+1}{\bm{d}}_{t}=\nabla J({\bm{x}}_{t+1})-\nabla J({\bm{x}}_{t}).$
From Eq. 8 we get the secant equation:
${\bm{B}}_{t+1}{\color[rgb]{0.171875,0.62890625,0.171875}\underbrace{\text{$({\bm{x}}_{t+1}-{\bm{x}}_{t})$}}_{\text{${\bm{s}}_{t}$}}}={\color[rgb]{0,0.265625,0.66796875}\underbrace{\text{$\nabla J({\bm{x}}_{t+1})-\nabla J({\bm{x}}_{t})$}}_{\text{${\bm{z}}_{t}$}}}\rightarrow{\bm{B}}_{t+1}{\color[rgb]{0.171875,0.62890625,0.171875}\text{${\bm{s}}_{t}$}}={\color[rgb]{0,0.265625,0.66796875}\text{${\bm{z}}_{t}$}}.$
To avoid explicitly computing the inverse matrix ${\bm{B}}_{t}^{-1}$, we can introduce an approximation ${\bm{H}}_{t}={\bm{B}}_{t}^{-1}$ and optimize it as follows:
${\bm{H}}_{t+1}=\underset{{\bm{H}}}{\operatorname{arg}\,\operatorname{min}}$ $\left\lVert{\bm{H}}-{\bm{H}}_{t}\right\rVert^{2}_{\bm{W}}$ $\rhd{\bm{H}}_{t+1}\ \text{close to}\ {\bm{H}}_{t}$ (10)
[…]
${\bm{H}}_{t+1}=({\bm{I}}-\rho_{t}{\bm{s}}_{t}{\bm{z}}_{t}^{\mathsf{T}}){\bm{H}}_{t}({\bm{I}}-\rho_{t}{\bm{z}}_{t}{\bm{s}}_{t}^{\mathsf{T}})+\rho_{t}{\bm{s}}_{t}{\bm{s}}_{t}^{\mathsf{T}},$ (11)
where $\rho_{t}=\frac{1}{{\bm{z}}_{t}^{\mathsf{T}}{\bm{s}}_{t}}$, ${\bm{s}}_{t}={\bm{x}}_{t+1}-{\bm{x}}_{t}$, and ${\bm{z}}_{t}=\nabla J({\bm{x}}_{t+1})-\nabla J({\bm{x}}_{t})$. This update equation allows us to iteratively refine the approximation ${\bm{H}}_{t}$ based on the current gradient information and the changes in the solution.
Figure 9: Inverse Hessian matrix approximation algorithm. Verification of requirements over the iterations. Left: Objective function value in Eq. 10; Right: Symmetry index of the inverse Hessian approximation refer to Eq. 12.
Figure 10: Visual comparison on AAPM. From top to bottom: the results under the following conditions: first $(n_{v}=32,N_{1})$, second $(n_{v}=64,N_{1})$, third $(n_{v}=128,N_{1})$. The display window is set to $\left[-1000,800\right]$ HU for the first two rows and to $\left[-200,300\right]$ HU for the last row.
A.2 Validation of adherence to BFGS
We validate the adherence of our method to the BFGS requirements. To achieve this, we present the constraint values of the optimization algorithm given in Eq. 10 using a test set image from AAPM, as illustrated in Fig. 9. The symmetry index is defined as follows:
$SI=\frac{1}{{n\cdot(n-1)}}\sum_{i=1}^{n}\sum_{j=1,j\neq i}^{n}\lvert{\bm{A}}_{ij}-{\bm{A}}_{ji}\rvert.$ (12)
Our results demonstrate the effectiveness of our approach in satisfying the essential conditions required by the BFGS algorithm. Notably, the symmetry index is consistently close to zero, indicating the symmetry of the matrix ${\bm{H}}_{t}$ at each iteration, which is the first constraint of the BFGS method. Furthermore, with regard to the objective function value, it is evident that it is close to zero, except for the initial approximation. This deviation can be attributed to the use of the identity matrix as the starting point.
Inverse Hessian matrix approximation visualizations.
Figure 7 depicts ${\bm{H}}_{t}$ at different iterations. These visualizations confirm the required symmetry of the matrix in each iteration. Additionally, the matrix ${\bm{H}}_{t}$ is close to the identity matrix at the second iteration, becoming more structured in the third iteration. This behavior aligns with expectations, as the matrix ${\bm{H}}_{t}$ is initialized as the identity matrix and updated based on gradient information and solution changes.
Visualization of reshaped rows.
To further understand the inverse Hessian matrix approximation structure, we depict the reshaped ($64\times 64$) first $40$ rows of ${\bm{H}}_{t}$ at iterations $9$ and $14$ in Fig. 8. These rows store gradient attention information used for updating the solution, consistent with the matrix ${\bm{H}}_{t}$ being updated based on gradient information and solution changes. In future work, we plan to explore the impact of ${\bm{H}}_{t}$ on the optimization process and its influence on reconstruction performance.
Appendix B More Ablation Study and visualization
Figure 11: Visual comparison on DeepLesion. From top to bottom: the results under the following conditions: first $(n_{v}=32,N_{1})$, second $(n_{v}=64,N_{1})$, third $(n_{v}=128,N_{1})$. The display window is set to $\left[-1000,800\right]$ HU for the first two rows and to $\left[-200,300\right]$ HU for the last row.
B.1 Ablation on Incept-Mixer
We further investigate the impact of the hyperparameters of Incept-Mixer on the reconstruction performance. We vary the patch size $p$ and the number of stacked Mixer layers $N$ and report the results in Tab. 7(a), and Tab. 7(b) respectively.
Impact of the path size.
We observe that increasing the patch size $p$ from $2$ to $4$ improves the performance ($+1.28$ dB and $+1.04$%) while further increasing the patch size from $4$ to $8$ decreases the performance ($-1.32$ dB and $-0.79$%).
We attribute this observed pattern to the trade-off between local and global features in the reconstruction process. When the patch size is small, such as $p=2$, the model focuses on capturing fine-grained local details, which can enhance reconstruction accuracy. As the patch size increases to $p=4$, the network gains a broader perspective by considering larger regions, leading to an improvement in performance. However, when the patch size becomes too large, for example, $p=8$, the model might start incorporating more global context at the expense of losing finer details. This can result in a decrease in performance as the model becomes less sensitive to localized patterns.
$p$ PSNR $\uparrow$ SSIM $\uparrow$
2 38.22 95.07
4 39.51 96.11
8 38.19 95.32
$N$ PSNR $\uparrow$ SSIM $\uparrow$
1 37.64 94.79
2 39.51 96.11
3 38.47 95.51
4 38.17 95.40
Table 7: Ablation of Incept-Mixer. (a) $p$ is the patch size; (b) $N$ is the number of stacked Mixer layers. The best performance is attained using $p=4$ and $N=2$.
Impact of the number of stacked Mixer layer.
We observe that increasing the number of stack $N$ from $1$ to $2$ improves the performance ($+1.86$ dB and $+1.31$%), while further increasing the patch size from $2$ to $3$ decreases the performance ($-1.03$ dB and $-0.60$%) and from $3$ to $4$ decreases the performance ($-0.30$ dB and $-0.11$%).
Similarly, when varying the number of stacked Mixer layers $N$, we observe a trend where an increase in $N$ initially contributes to improved performance, as the model can capture more complex features and relationships. However, as $N$ continues to grow, the network may encounter diminishing returns, and the benefits of additional layers diminish, potentially leading to overfitting or increased computational overhead.
Robustness to hyperparameters.
Hence, there exists an optimal trade-off between the patch size $p$ and the number of stacked Mixer layers $N$, but the model performs similarly for a wide range of values. In our experiments, we use $p=4$ and $N=2$ for all the datasets, which highlights the robustness of our method to these hyperparameters.
B.2 More visualization results
Fig. 10 displays supplementary visualizations of our approach on AAPM. Our method consistently produces high-quality reconstructions across all views. Notably, among state-of-the-art techniques, QN-Mixer excels in reconstructing fine-grained details. For instance, it accurately captures small vessels in the first row, delicate soft tissue structures in the second row, and sharp boundaries in the third row.
In Fig. 11, we showcase additional visualizations of our method applied to DeepLesion. QN-Mixer demonstrates superior performance across all views, yielding high-quality reconstructions. This is particularly evident in the challenging scenario of $32$ views, where our method outperforms others in capturing fine-grained details, such as small vessels and lesions. Importantly, these results are achieved with fewer iterations compared to alternative unrolling networks like RegFormer.
B.3 Iterative results visualization
Figure 12: QN-Mixer’s intermediate reconstructions using AAPM with $(n_{v}=32,N_{1})$. Display window is set to $\left[-1000,800\right]$ HU.
In order to demonstrate the effectiveness of QN-Mixer, we present a series of intermediate reconstruction results in Fig. 12. These results illustrate the progression of the reconstruction process at different iterations of our method. By examining the reconstructed outputs at each iteration, our goal is to offer insights into the evolution of image quality. Notably, we observe that the improvement in quality, as quantified by the PSNR and SSIM values of each iteration, does not consistently increase with each iteration (see Iteration $10$ in Fig. 12). We suspect that the observed unexpected behavior may arise from the variation of the objective function (i.e. Eq. 2) around the point $t$ in the unrolled network, which is dependent on a learnable gradient regularization term Incept-Mixer.
B.4 Reconstruction error visualization
Figure 13: Reconstruction errors with LEARN, RegFormer, and QN-Mixer using $(n_{v}=32,N_{0})$. Images are ordered left to right by SSIM, with the first column showing the worst reconstruction among $214$ patient images. The second and third columns represent the $1/3$ and $2/3$ percentiles, respectively, and the last column corresponds to the best reconstruction with the highest SSIM.
We present the reconstruction error of QN-Mixer in comparison to LEARN and RegFormer in Fig. 13. The images are organized from left to right based on the SSIM value. As illustrated, our method consistently produces high-quality reconstructions. In the most challenging scenario ($n_{v}=32$) with no added noise, and for the least favorable image, our method achieves a reconstruction with an SSIM of $93.53\%$, maintaining notably satisfactory performance compared to RegFormer with an SSIM of $91.30\%$. For the best reconstruction across all methods, our method achieves an SSIM of $97.72\%$, while RegFormer achieves an SSIM of $97.48\%$. These results demonstrate the robustness of our method when dealing with challenging scenarios.
Method $n_{v}=32$ $n_{v}=64$ $n_{v}=128$
SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$
FBP 72.88 18.97 83.42 22.13 91.75 24.85
FBPConvNet [19] 63.91 20.94 73.02 24.12 80.60 25.74
DuDoTrans [49] 60.51 19.09 79.75 25.00 85.65 27.23
Learned PD [1] 67.99 21.92 83.79 25.51 85.42 25.86
LEARN [7] 79.70 24.46 84.44 26.74 88.16 26.20
RegFormer [54] 72.45 23.69 77.33 25.46 84.99 28.22
QN-Mixer (ours) 86.17 25.95 94.56 30.95 97.04 33.98
Table 8: Quantitative results of the reconstruction of the cropped OOD circle. Bold: Best, under: second best.
Appendix C More experiments
C.1 Out-of-Distribution
(a) $n_{v}=32,N_{0}$
(b) $n_{v}=64,N_{0}$
(c) $n_{v}=128,N_{0}$
Figure 14: Visualization of 5 samples of the OOD circle texture reconstruction. From each figure and from left to right, we show the ground truth, FBP, FBPConvNet, DuDoTrans, Learned PD, LEARN, RegFormer and QN-Mixer.
OOD circle performance across entire image  In the main text, we assess the robustness of methods to a simple out-of-distribution scenario, where an unseen during training, white circle is inserted, and computing SSIM and PSNR metrics for the entire image. We achieved the best performance across all views in this setting (Tab. 3). We note, however that this setting mixes the inherent performance on clean data and its ability to handle unseen patterns.
OOD circle performance in anomalous region  To further isolate the capacity of models to reconstruct unseen OOD patterns (i.e., white circle), we extend our evaluation. Instead of evaluating the whole image, we compute the reconstruction performance of a crop region containing the white circle, thus isolating the reconstruction performance exclusively to the circle region.
For that, we randomly selected 5 samples from the AAPM test set depicted. The evaluated data is depicted in Fig. 14, and the overall performance across the complete set of 214 patient images is summarized in Tab. 8 for 3 views.
Our method significantly outperforms the second-best across all views in both SSIM and PSNR. For the most challenging case of $32$ views, we surpass the second best by $+6.47\%$ and $+1.49$ dB. With $64$ views, our performance exceeds the second best by $+10.12\%$ and $+4.21$ dB. In the easiest case of $128$ views, we outperform the second best by $+5.29\%$ and $+5.76$ dB. As anticipated, all methods exhibit degraded performance when focusing on the circle region, and the gap between our method and the second-best widens compared to the complete image. Moreover, the numerical results in Tab. 8 align with the visualizations in Fig. 14.
Train AAPM $n_{v}{=}32$ AAPM $n_{v}{=}128$
Test DeepLesion $n_{v}{=}32$ AAPM $n_{v}{=}32$ AAPM $n_{v}{=}64$
Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
DuDoTrans 28.69 73.77 25.60 51.01 32.38 82.35
LEARN 37.21 93.46 28.20 66.10 34.40 87.21
RegFormer 37.97 94.01 29.83 72.53 37.78 92.64
QN-Mixer (ours) 38.28 94.66 30.43 80.03 37.16 91.81
Table 9: Quantitative results of the reconstruction performance of OOD cases: anatomy and geometry Bold: Best, under: second best.
Anatomy and Geometry OOD.  We explore more complex out-of-distribution (OOD) scenarios, particularly focusing on changes in anatomy and geometry. For the geometry aspect, we train our methods using AAPM data and evaluate them on DeepLesion, maintaining a fixed geometry parameter of $n_{v}{=}32$. Concerning anatomy, we test under the most challenging scenario, which involves training on AAPM data with $n_{v}{=}128$ and subsequently testing on AAPM data with two lower geometries: $n_{v}{=}32$ and $n_{v}{=}64$. The outcomes of this experiment are detailed in Tab. 9.
  • anatomy: we outperform RegFormer by $+0.31$ dB and $+0.65\%$.
  • geometry $n_{v}{=}32$: we outperform Regformer by $+0.60$ dB and $+7.50\%$ in most challenging setting.
  • geometry $n_{v}{=}64$: we lag behind Regformer by $-0.62$ dB and $-0.83\%$ in the easiest setting.
C.2 Results when A is the $90^{\circ}$ limited view CT
To further evaluate the robustness of our method, we conducted experiments on a more challenging CT inverse problem, specifically involving a $90^{\circ}$ limited-angle CT setup. In this scenario, the system matrix $A$ comprises a restricted number of views, resulting in significant artifacts and distortions in the reconstructed images. We compared the performance of our proposed QN-Mixer, with that of RegFormer on this demanding task. Our method demonstrated superior performance, achieving promising results with a PSNR of $30.17$ dB and a SSIM of $90.01\%$. This outperformed RegFormer by $+1.03$ dB in PSNR and $+0.45\%$ in SSIM. Visual results illustrating the effectiveness of our approach are provided in Fig. 15.
Figure 15: Visual comparaison on AAPM with Limited-angle CT. The system matrix $A$ is a $90^{\circ}$ limited view. The display window is set to $\left[-1000,800\right]$ HU.
C.3 MBIR evaluation
To demonstrate the superiority of our method, we conducted an additional comparison with the state-of-the-art model-based iterative reconstruction (MBIR) technique, which is widely employed in clinical settings. Our evaluation involved testing our approach against MBIR using datasets with varying numbers of views, specifically $n_{v}{=}32,64$, and $128$, with low noise added to sinograms. The corresponding PSNR values obtained with MBIR were $22.94$, $28.80$, and $34.33$ dB, with SSIM values of $67.70\%$, $72.60\%$, and $80.68\%$, respectively.
Our method, QN-Mixer, consistently outperformed MBIR across all views, showcasing an average improvement of $+12.72$ dB in PSNR and $+23.25\%$ in SSIM. These results underscore the robustness and effectiveness of our approach compared to MBIR, a benchmark technique widely utilized in clinical practice.
C.4 Noise Power Spectrum analysis
(a) $n_{v}=32,N_{1}$
(b) $n_{v}=64,N_{1}$
(c) $n_{v}=128,N_{1}$
Figure 16: Noise Power Spectrum (NPS) Analysis in comparison to state-of-the-art methods. The x-axis represents normalized frequency in cycles per pixel ($\text{px}^{-1}$), and the y-axis represents noise power spectrum ($\text{HU}^{2}\text{px}^{2}$). Display windows are configured as $\left[-1000,800\right]$ HU. Mean and peak frequencies are intricately linked to noise texture, with finer textures correlating to higher mean and peak frequencies in the NPS. Our method exhibits the highest peak frequencies, indicating that our reconstructed images feature the most refined noise texture among all compared methods.
Figure 17: ROIs for NPS Analysis: Red squares denote $20\times 20$ pixel ROIs distributed evenly across two circular regions. The first circle (radius $25$) holds $8$ ROIs, and the second circle (radius $50$) has $20$ ROIs. Both circles, centered at the image center, include a total of $29$ ROIs per image. This standard positioning in the CT community underscores the clinical diagnostic importance of the image center.
We conducted a comprehensive examination of the noise characteristics in our reconstructed images through noise power spectrum (NPS) analysis. NPS serves as a metric, quantifying the magnitude and spatial correlation of noise properties, or textures, within an image. It is derived from the Fourier transform of the spatial autocorrelation function of a zero-mean noise image.
NPS analysis was performed on a configuration of Regions of Interest (ROIs) as depicted in Fig. 17. This process was applied to all $214$ images from the AAPM test set and for three different views ($32$, $64$, and $128$). The average 1D curves were generated by radially averaging the 2D NPS maps, and the results are presented in Fig. 16.
The area under the NPS curve is equal to the square of the noise magnitude. Importantly, the ordering of methods based on noise magnitude corresponds to the ranking observed in our quantitative experiments for PSNR and SSIM in the main text. For example, FBP, which exhibits the lowest noise magnitude, also performs the poorest in terms of PSNR and SSIM. Conversely, our method, with the highest noise magnitude, stands out as the top performer in both PSNR and SSIM metrics. Furthermore, the mean and peak frequencies serve as key indicators of noise texture or “noise grain size”, where higher frequencies denote finer texture. Remarkably, our method showcases superior mean and peak frequencies compared to other methods, suggesting a finer noise texture or smaller grain size.
This alignment with good clinical practice standards reinforces the robust performance of our method in capturing and preserving image details, as supported by both quantitative metrics and noise analysis.
Appendix D Reproducibility
1class GradientFunction(nn.Module):
2 def __init__(self, regularizer):
3 self.regularizer = regularizer
4 self.lambda = nn.Parameter(torch.zeros(1))
5
6 def forward(self, physics, y, x):
7 y_t = physics.forward_operator(x)
8 # Compute the regularization term
9 reg_x = self.regularizer(x)
10 # Compute the data fidelity term
11 y_dft = y_t - y
12 # Compute the backprojection
13 x_dft = physics.backward_operator(y_dft)
14 g = self.lambda * x_dft + reg_x
15 return g
16
17class QN_Iteration(nn.Module):
18 def __init__(self, gradient_function):
19 self.gradient = gradient_function
20
21 def latent_bfgs (self, h, s_t, z_t):
22 I = torch.eye(len(s_t))
23 rho_t = 1. / torch.dot(z_t, s_t)
24 u_t = I - torch.outer(s_t, z_t) * rho_t
25 d_t = I - torch.outer(z_t, s_t) * rho_t
26 return (torch.matmul(u_t, torch.matmul(h, d_t))
27 + (torch.outer(s_t, s_t) * rho_t))
28
29 def forward(self, physics, encoder, decoder,
30 y, x, h, r, is_last):
31 # Compute latent direction s_t
32 s_t = -torch.matmul(h, r)
33 d = decoder(s_t)
34
35 # Update the reconstruction
36 x = x + d
37 # Return x if it is the last iteration
38 if is_last:
39 return x, h, r
40 else:
41 r_p = encoder(self.gradient(physics, y, x))
42 z_t = r_p - r
43 with torch.no_grad():
44 h = self.latent_bfgs(h, s_t, z_t)
45 return x, h, r_p
Algorithm 3 Minimal QN-Mixer pseudo-code
All our experiments are fully reproducible. While the complete algorithm is already provided in the main paper (see Algorithm 2), we additionally present a PyTorch pseudo-code for enhanced reproducibility in Sec. D.1. We furnish comprehensive references to all external libraries used in Sec. D.2. Detailed information regarding the initialization of our model can be found in Sec. D.3. The precise parameters of our regularizer, Incept-Mixer architecture, are available in Sec. D.4. We outline the exact data splits utilized across the paper for the AAPM dataset in Sec. D.5. Lastly, to facilitate the reproduction of our out-of-distribution (OOD) protocol, we provide the pseudo-code in Sec. D.6.
D.1 QN-Mixer pseudo-code
Our QN-Mixer algorithm is introduced in Algorithm 2, and for improved reproducibility, we present a PyTorch pseudo-code in Algorithm 3. The fundamental concept underlying unrolling networks lies in having a modular gradient function, denoted as $\nabla J({\bm{x}}_{t})$, which can be easily adapted to incorporate various regularization terms. Subsequently, the core element is the unrolling iteration block responsible for updating both the solution ${\bm{x}}_{t}$ and the inverse Hessian approximation ${\bm{H}}_{t}$. The update of the inverse Hessian approximation is executed through the latent BFGS algorithm. Notably, each iteration call takes the physics operator as input, tasked with computing the forward and pseudo-inverse operators for the CT reconstruction problem, along with the gradient encoder and direction decoder, which are shared across all iterations. For a more in-depth understanding, refer to Algorithm 3. Note the employment of torch.no_grad() to inhibit the computation of gradients for the inverse Hessian approximation. Since there is no necessity to compute gradients for this variable, given that it is updated through the latent BFGS algorithm.
Within these two modules, second-order quasi-Newton methods can be seamlessly incorporated by simply modifying the latent BFGS algorithm or the regularization term, offering flexibility to the user.
D.2 External libraries used
We utilized the following external libraries to implement our framework and conduct our experiments:
D.3 QN-Mixer’s parameters initialization
To enhance reproducibility, we provide the parameters initialization of QN-Mixer. First, for the gradient function, we initialize the CNNs of Incept-Mixer using the Xavier uniform initialization. The multi-layer perceptron of the MLP-Mixer is initialized with values drawn from a truncated normal distribution with a standard deviation of $0.02$. The $\lambda_{t}$ values are initialized to zero, and the inverse Hessian approximation ${\bm{H}}_{0}$ is initialized with the identity matrix ${\bm{I}}$. Second, for the latent BFGS, both the encoder and decoder CNNs are initialized with the Xavier uniform initialization.
D.4 Incept-Mixer’s architecture
For enhanced reproducibility, we present the architecture of Incept-Mixer in Tab. 11. The Incept-Mixer architecture consists of a sequence of Inception blocks, followed by Mixer blocks. Each Mixer block comprises a channel-mixing MLP and a spatial-mixing MLP. The MLPs are constructed with a fully-connected layer, a GELU activation function, and another fully-connected layer. Ultimately, the regularization value is projected to the same dimension as the input image through a patch expansion layer, which is composed of a fully-connected layer and a CNN layer.
Patient ID L067 L109 L143 L192 L286 L291 L096 L506 L333 L310
#slices 224 128 234 240 210 343 330 211 244 214
Training
Validation
Testing
Table 10: AAPM dataset split specification. The validation set comprises images from patient L333, and testing utilizes images from patient L310. The images from the remaining patients have been designated for training purposes.
Stage Layers #Param (k) Output size
Input - - $1\times 256\times 256$
InceptionBlock-1
convblock1:
conv1-1: K1C16S1P0
prelu1-1
convblock2:
conv2-1: K1C16S1P0
prelu2-1
conv2-2: K3C32S1P1
prelu2-2
convblock3:
conv3-1: K1C16S1P0
prelu3-1
conv3-2: K5C32S1P2
prelu3-2
convblock4:
maxpool4-1: K3S1P1
conv4-1: K1C16S1P0
prelu4-1
17.6 $96\times 256\times 256$
PatchEmbed-2
conv2-1: K4C96S4P0
rearrange2-1: bchw $\rightarrow$ bhwc
$145.5$ $96\times 64\times 64$
MixerLayer-3 ×$\times(N=2)$
layernorm3-1: D96
rearrange3-1: bhwc $\rightarrow$ bcwh
heightmlp3-1:
linear3-1: D64O256
gelu3-1
linear3-2: D256O64
rearrange3-2: bcwh $\rightarrow$ bchw
widthmlp3-1:
linear3-3: D64O256
gelu3-2
linear3-4: D256O64
rearrange3-3: bchw $\rightarrow$ bhwc
layernorm3-2: D96
channelmlp3-1:
linear3-5: D96O384
gelu3-3
linear3-6: D384O96
$140.8\times 2$ $96\times 64\times 64$
PatchExpand-4
linear4-1: D96O1536
layernorm4-1: D96
conv4-1: K1C1S1P0
$47.71$×256×25612562561$\times$ 256$\times$ 2561 × 256 × 256
Table 11: Incept-Mixer architecture. K-C-S-P represents the kernel, channel, stride, and padding configuration of CNNs, while D-O indicates the input and output dimensions of linear layers.
D.5 AAPM dataset splits
In our experiments, we use the AAPM 2016 Clinic Low Dose CT Grand Challenge public dataset [35], which holds substantial recognition as it was formally established and authorized by the esteemed Mayo Clinic. To ensure the integrity of our evaluation process, we followed the precedent set by [7, 54] and created the training set using data from eight patients, while reserving a separate patient for the testing and validation sets. This approach guarantees that no identity information is leaked during test time. Our specification is presented in Tab. 10.
D.6 Robustness eval. protocol for OOD scenarios
1def add_circle_ood(img, value=1):
2 h, w = img.shape[::-1][:2]
3 radius = np.random.randint(5, 20)
4 c_x = np.random.randint(radius, w-radius)
5 c_y = np.random.randint(radius, h-radius)
6 center = (c_x, c_y)
7
8 Y, X = np.ogrid[:h, :w]
9 dist_x = (X - center[0])**2
10 dist_y = (Y - center[1])**2
11 dist_from_center = np.sqrt(dist_x + dist_y)
12 mask = dist_from_center <= radius
13 img[0, mask] = value
14 return img
Algorithm 4 add_circle_ood pseudo-code.
In medical imaging, it’s crucial to develop methods that generalize to scans with lesions or anomalies, and assessing the model’s capability to reconstruct abnormal data holds significant relevance, as test patient data may deviate from the training data in clinical applications. To this end, we design a simple protocol specifically crafted for evaluating the effectiveness of methods when handling abnormal data. In this case, a white circle mimicking an out-of-distribution texture, which was never seen during training, is forged into CT images with noise-free sinograms. The pseudo-code to realize this is provided in Algorithm 4. Then, performance can be evaluated on the entire image as shown in Tab. 3, or on a cropped region within the circle as detailed in Tab. 8.
We strongly advocate for future research endeavors to embrace and employ this protocol as a standard for evaluating the robustness of reconstruction methods.
Appendix E Limitations
Our approach inherits similar limitations from prior methods [7, 54]. First, our method entails a prolonged optimization time, stemming from the utilization of unrolling reconstruction networks [7, 54], in contrast to post-processing-based denoising methods [19, 49]. While our method represents the fastest unrolling network, there is still a need to address the existing gap. Integrating Limited-memory BFGS into our QN-Mixer framework is an interesting research direction for accelerating training. Second, while we have assessed our method using the well known AAPM low-dose and DeepLesion datasets and compared it with several state-of-the-art methods, the evaluation is conducted on images representing specific anatomical regions (thoracic and abdominal images). The generalizability of our method to a broader range of datasets, which may exhibit diverse characteristics or variations, remains unclear. Third, the acquisition of paired data has always been an important concern in clinic. Combining our approach with unsupervised training framework to overcome this limitation can be an exciting research direction. Finally, the incorporation of actual patient data into our training datasets raises valid privacy concerns. Although the datasets we utilized underwent thorough anonymization and are publicly accessible, exploring a solution that can effectively operate with synthetic data emerges as an intriguing avenue to address this challenge.
Appendix F Notations
We offer a reference lookup table, available in Table 12, containing notations and their corresponding shapes as discussed in this paper.
Notation Shape Value(s) Description
$n_{v}$ $\mathbb{N}^{*}$$\{32,64,128\}$ The number of projection views
$n_{d}$ $\mathbb{N}^{*}$ 512 The number of projection detectors
$h$ $\mathbb{N}^{*}$ $256$ Height of the image
$w$ $\mathbb{N}^{*}$ $256$ Width of the image
$c$ $\mathbb{N}^{*}$ $1$ Channels of the image
$l_{h}$ $\mathbb{N}^{*}$ $64$ Latent height
$l_{w}$ $\mathbb{N}^{*}$ $64$ Latent width
$m=n_{v}\times n_{d}$ $\mathbb{N}^{*}$ $n_{v}\times 512$ Data (sinogram) size
$n=h\times w$ $\mathbb{N}^{*}$ $256\times 256$ Image size
[…]
$T$ $\mathbb{N}^{*}$ 14 Number of iterations of our method
[…]
${\bm{y}}$ $n_{v}\times n_{d}$ Sparse sinogram
${\bm{A}}$ $\mathbb{R}^{n\times m}$ The forward model (i.e. discrete Radon transform)
${\bm{A}}^{\dagger}$ $\mathbb{R}^{m\times n}$ The pseudo-inverse of ${\bm{A}}$
${\bm{x}}_{0}$ $\mathbb{R}^{h\times w\times c}$ ${\bm{A}}^{\dagger}{\bm{y}}$ Initial reconstruction
[…]
${\bm{x}}_{t}$ $\mathbb{R}^{h\times w\times c}$ Reconstructed image at iteration $t$
$\nabla_{\bm{x}}J({\bm{x}}_{t})$ $\mathbb{R}^{h\times w\times c}$ Gradient value at iteration $t$
${\bm{H}}_{t}$ $\mathbb{R}^{(l_{h}\cdot l_{w})\times(l_{h}\cdot l_{w})}$ Approximation of the inverse Hessian matrix at iteration $t$
${\bm{I}}^{n\times n}$ $\mathbb{N}^{n\times n}$ Identity matrix of size $n\times n$
${\bm{f}}_{t}$ $\mathbb{R}^{h\times w\times d}$ Feature map after the Inception block at iteration $t$
${\bm{e}}_{t}$ $\mathbb{R}^{\frac{h}{p}\times\frac{w}{p}\times d}$ MLP-Mixer embeddings
[…]
$p$ $\mathbb{N}^{*}$ $4$ Stride and kernel size in the patchification Conv 2D net
$N$ $\mathbb{N}^{*}$ $2$ Number of stacked Mixer layers
${\mathcal{G}}(\cdot)$ Learned gradient of the regularization term (i.e. the Incept-Mixer model)
${\mathcal{G}}({\bm{x}}_{t})$ $\mathbb{R}^{h\times w\times c}$ Regularization term at step $t$
$\mathcal{E}(\cdot)$ The gradient encoder
$\mathcal{D}(\cdot)$ The direction decoder
$k$ $\mathbb{N}^{*}$$\{2,3,4,5\}$ Number of Downsampling stacks in the encoder
${\bm{f}}_{\mathcal{E}}=2^{k}$ $\mathbb{N}^{*}$$\{4,8,16,32\}$ Downsampling factor of the gradient in the encoder
$w_{l}={w}/{{\bm{f}}_{\mathcal{E}}}$ $\mathbb{N}^{*}$$\{64,32,16,8\}$ Number of columns of the down-sampled gradient
$h_{l}={h}/{{\bm{f}}_{\mathcal{E}}}$ $\mathbb{N}^{*}$$\{64,32,16,8\}$ Number of rows of the down-sampled gradient
${\bm{r}}_{t}=\mathcal{E}(\nabla_{{\bm{x}}}J({\bm{x}}_{t}))$ $\mathbb{R}^{l_{h}\cdot l_{w}}$ Latent representation of the gradient
$s_{t}=-{\bm{H}}_{t}{\bm{r}}_{t}$ $\mathbb{R}^{l_{h}\cdot l_{w}}$ Direction in the latent space
$\rho_{t}={({\bm{z}}_{t}^{\mathsf{T}}{\bm{s}}_{t})}^{-1}$ $\mathbb{R}^{l_{h}\cdot l_{w}}$ BFGS divider variable
$N_{0}$ Zero noise added to the sinogram
$N_{1}$ 5% Gaussian noise, $1\times 10^{6}$ intensity Poisson noise
$N_{2}$ 5% Gaussian noise, $5\times 10^{5}$ intensity Poisson noise
Table 12: Lookup table of notations and hyperparameters used in the paper.